Astronomy Report – January 27th



Name: Rob Hunt
Crew: 291
Date: 27/01/2024
Solar Features Observed: Sun spots
Images submitted with this report: None
Time in observatory: 10:00 to 16:30 (less 30 minutes lunch break in HAB). Optimum period for observing is approx. 10:00 – 15:00
Problems Encountered:
Scope was found pointing North.
Interior hook for padlock as shown in Guide, is not there.
Broken eye-piece tightening screw found in top drawer.
Shutter Control box operates very poorly. Initially, lower shutter would not open at all but eventually responded. Order in which shutters open changes randomly or they don’t both open. Power cycling might make a difference but must be ‘Off’ for 10-15 minutes. Sometimes lower shutter stops at about ¼ open. There is a clunk/stall in the worm drive about ⅓ way open but iot doesn’t stop the movement. Worm thread is not lubricated.
Lower shutter fits over its seating rims very unevenly leaving gaps on the sides and at the overlap with upper shutter when closed. Lower shutter has four loose screws on outside of the edge that closes against the upper shutter. The hinges attaching iot to the dome are stiff (with paint?), not secured well and wobble.
Upper shutter follows its guides OK but is very noisy and unlubricated. When closing against the lower shutter, it leaves large gaps – caused by misaligned lower shutter.
Azimuth motor works OK but dome wheels don’t always run on the track. Dome and walls do not match up around the track and wheels hit gaps at mismatched wall segments. Nylon drive ring guide above controller shunts back and forth with loud clunking sound and makes dome shudder.
Azimuth drive control paddle has a light metal label immediately under the toggle switch. It has “On” in red on opposite sides buts is misaligned relative to the “Left” and “Right” labels on the outside of the paddle.
Guide book pages came away from spine and fell out.
In Guide book, “Alignment Success” is not displayed. Instead “Select Alignment” appears. I selected “Solar System Alignment” then “Sun”. I then got “Don’t look at Sun” warning.
All alignment problems (p 13 Guide) were experienced.
No instructions in Guide to remove lens cap.
No instructions in Guide to remove red plastic eyepiece socket plug.
One eyepiece holding screw is seized in position.
One eyepiece holding screw is snapped off leaving thread in the collar. Screw head found in drawer is a match for the other half.
Eyepiece cannot be securely inserted. It wobbles when used in situ.
ZWO cable was found dangling on the floor.
ZWO camera sits in eyepiece collar very loosely because of missing tightening screws.
Locating “USB Extension” not described in the Guide.
SharpCap update was downloade but not installed because password protected (v 4.0.9538)
No “Gamma” controls found in SharpCap menus.
Closing Proceedure with scope Controller – Menu doesn’t say “Goto Home Position” but says “Go to”.
There is no “Hibernate” in “Home Position”
Today, the “Hibernation” position was very different to where scope was found.
Button #7 on Control paddle goes Menu >> Utilities >> Home Position >> Goto

Astronomy Report – January 23rd



Name: Rob Hunt
Crew: 291
Date: 23/01/2024
Solar Features Observed: None
Images submitted with this report: None
Problems Encountered: Total cloud 100% all day. Seeing zero. Overnight the laptop was fully re-charged from Critical (4%)

GreenHab Report – January 14th

Crew 290 GreenHab Report 14-01-2024 (Sol 7)

Greenhab Officer: Ben Kazimer

Heater cycled throughout day due to warm weather but is currently on, fan off, door closed

Average temperatures: 0925: 77.4⁰F, 25%, 1648: 63.5⁰F, 28%

Hours of supplemental light: 4.5 (4 overnight and ½ from watering after sunset)

Daily water usage for crops: 7 gallons

Daily water usage for research and/or other purposes: None!

Water in Blue Tank 151.05 gallons

Time(s) of watering for crops:

0925: 5 gallons watering, ½ gallon misting, 1257: ¼ gallon misting, 1648: ¼ gallon misting, 1833: 1 gallon watering

Changes to crops: rearranged tomatoes and cucumbers for more separation on main planter table, also trimmed their leaves


Woahhhhhhh we’re halfway there! While the plants did not listen to any Bon Jovi this sol, Crew 290 is halfway through our rotation. The plants seemed a bit melancholic with stressed leaves, but the fruit looked great so I’m chalking that up to yestersols dosage of Rat Pack (+ Louis Prima) vocals. To remedy the leaves, the cucumbers and tomatoes did another Martian shuffle today, this time to increase separation between pots as branches and leaves were getting a little too close for everyone’s comfort. This did come with the price of my observation bench as I needed more planter real estate, but I am willing to make such sacrifices for the increased health of these plants. During the afternoon check-in, I noticed Tomato No.4 is really excelling after previous struggles! They are the Sol 7 rising star, winning the title “Comeback Kid”. I also made sure all the herbs were staying in their designated growing regions and did my best to reduce any physical altercations. Finally, I added a little compost to Riya’s spinach, mint, and onion to encourage growth (the onions are doing great!).

Harvest: None :/

Support/supplies needed: None!

Operations Report – January 14th

Crew 290 Operations Report 14-01-2023

SOL: 7

Name of person filing report: Anna Tretiakova

Non-nominal systems: None

Notes on non-nominal systems: Not Applicable


Spirit rover used: No

Hours: (Before EVA)

Beginning Charge: (Before EVA)

Ending Charge: (After EVA and before recharging)

Currently Charging:

Opportunity rover used: No

Hours: (Before EVA)

Beginning Charge: (Before EVA)

Ending Charge: (After EVA and before recharging)

Currently Charging:

Curiosity rover used: No

Hours: (Before EVA)

Beginning Charge: (Before EVA)

Ending Charge: (After EVA and before recharging)

Currently Charging:

Perseverance rover used: No

Hours: (Before EVA)

Beginning Charge: (Before EVA)

Ending Charge: (After EVA and before recharging)

Currently Charging:

General Notes on Rovers: In nominal condition.

Summary of Hab Operations:

WATER USE: 34.5 gallons

Water (static tank): 348.7 gallons

Static tank pipe heater (on or off): on

Static tank heater (on or off): on

Toilet tank emptied: Yes

Summary of Internet: internet is nominal during times 18:00 to 22:00

Summary of Suits and Radios: Radios nominal. Suits nominal.

Summary of Greenhab Operations:

WATER USE: 7.0 gallons

Heater: on

Supplemental Light: off

Harvest: No

Summary of Science Dome Operations:

All collected soil samples were continued to be put into microtubular devices, imaged and incubated. The pH sensors were tested and prepared for tomorrow’s EVA.

Dual Split Heater or AC: On

Summary of RAM Operations: (Tools used, work done):

Measuring tape was used to measure the water level. The soldering iron and all the supplies for it were used successfully to fix the temperature sensor. Some assorted caps, screws, washers and nuts were taken to test the dexterity of crew 290 gloves. All materials were returned at the end of the day.

Summary of any observatory issues: Nominal

Summary of health and safety issues:

All crew in good health and happy spirits. No safety issues.

Questions, concerns and requests to Mission Support:

As mentioned above, the temperature sensor was fixed. Below I have attached images of how it currently looks inside. The 9V battery clip was opened up, the folded wires in it were taken out and soldered directly to the wires that come from the motherboard. Both wires were then covered in corresponding electrical tape. Although the battery can be changed, the future users should be mindful and careful. The temperature sensor is now nominal and ready for hardcore field work.

Trash from the bathroom was changed and put in the rear air lock for collection. That trash can be burned.

Sol Summary – January 14th

Crew 290 Sol Summary Report 14-01-2024

Sol: 7

Summary Title: pHun Day!

Author’s name: Madelyn Hoying

Mission Status: Nominal

Sol Activity Summary:

Our rest day started with crepes with peanut butter, Nutella, and strawberries courtesy of Freebee (Anja). Chopper (Rebecca), Funk (Ben), Roots (Anna), and Freebee spent some time loading microfluidic devices with samples collected yesterday, while Melon (Madelyn) evaluated field photos and rock samples to characterize the sample sites. After a lunch of Funk’s fried rice, Nicole (PODO) and Chopper continued pH measurements while Roots evaluated glove dexterity with Freebee and Melon. The crew is enjoying some personal time and getting ready for our team bonding activities tonight.

Look Ahead Plan: Sample collection and analysis

Anomalies in work: none

Weather: Warm and windy

Crew Physical Status: nominal

EVA: none

Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, EVA Request

Support Requested:

Trash pickup in the rear airlock.

Supplemental Operations Report – January 14th

Date: 1/14/2024
Name of person filing report: Ben Stanley
Reason for Report: Routine
Non-Nominal Systems: Intern trailer heater, Perseverance rear view mirror, Perseverance parking brake, Crew car

Power system:
Solar: Nominal.
SOC Last 24 hours: Max 100%; Min 60%; Avg 74.3%.
VDC Last 24 hours: Max 59.1V; Min 40.15; Avg 49.00V.
Generator run time: 4683.1 hours. Oil and air filter changed on 01/12/2024

Propane Readings:
Station Tank: 77%
Director Tank: 82%
Intern Tank: 91%
Generator Tank: 77%

Hab Static Tank – 403.2 gallons
GreenHab – 158.1 gallons
Outpost tank – 400 gallons
Science Dome – 0 gallons
Septic Tanks emptied: Yes

Sojourner rover used: No.
Hours: 196.2
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: Yes
Notes on Rovers: Perseverance parking brake non-functional and rear-view mirror cracked.

ATV’s Used: None. Nothing to report.

Hab Car used and why, where: To Hanksville for supplies. Need to be serviced.
Crew Car used and why, where: N/A
General notes and comments: Crew car in Loa for repairs.

Summary of Internet: All nominal.

EVA suits and radios:
Suits: Blown fuse in suit #8 was successfully repaired by crew engineer.
Comms: All nominal

Campus wide inspection, if action taken, what and why: All nominal. Two fire extinguishers in GJ for servicing.

Summary of Hab Operations: UV lamp in water filter needs to be replaced. Pending.

Summary of GreenHab Operations: All nominal

Summary of SciDome Operations: All nominal

Summary of Observatories Operations: Adhesive for weather stripping in robotic observatory is in transit.

Summary of RAM Operations: All nominal

Summary of Outpost Operations: Time machine utilized to bring intern trailer heater back from the future (Dec 2025), still working on repairs. Fire extinguisher still in GJ for servicing.

Summary of Health and Safety Issues: All nominal.

Journalist Report – January 13th

Journalist Report

Nicole Chan, Crew 290 Journalist


Sol 6

Our crew awoke feeling rested for the long day ahead. We have two target sample selection sites for today’s EVA #7, to be led by Rebecca (Chopper), Ben (Funk), Anna (Roots), and myself (PODO). But before that, we need to prepare solutions to test the pH of the soil samples in the field. So, the crew went off to the ScienceDome to begin preparations. Madelyn (Melon) reviewed geological notes of the Candor Chasma region to give the EVA crew an idea of what a viable sample collection site could look like. Meanwhile, Chopper led Anja (Freebee) on loading the microfluidic devices with the samples from Sol 4. We found more pH meters! As Funk and I calibrated them, we realized that their functionality will make them suitable for the field, which made today’s EVA data even more fruitful.

After a lunch of pasta, the EVA crew suited up and drove the rovers to Candor Chasma. The road was bumpy but made for a fun (but also safe) ride. Upon reaching the area around Compass Rock, the crew decided to disembark and start searching for viable sites. We were looking for strips of darker red soil layered by white layers which denote gypsum deposits (where there could be signs of life!). After selecting a site, we got to work taking temperature, ATP, and pH measurements of the soil. We also took soil samples to see if there are any signs of microbial life too. As the clouds were rolling in and it was nearing sunset, we stepped up our measurements at our second site as well before heading back to the Hab. We returned on time and got to share our findings with the rest of the crew.

The crew has also put their heads together to flesh out the science protocols for processing these samples to best maintain consistency and get statistically significant results (with limited chemical supplies as well). We found great workarounds that still maintain the fidelity of our tests so we’ll be busying ourselves with those measurements and imaging in the coming days. Tomorrow, however, will be our rest day. After an Earth week of getting situated and gathering samples, it’s a well-deserved break for all of us. Stay tuned for tomorrow’s report to see what kind of fun things we’ll get up to!

Photo of the Day: pHenomenalogy studies.JPG

Crew 290 Mid-mission Research Report – 13Jan2024

[title Midmission Research Report – January 13th]
[category science-report]

Mars Desert Research Station

Mid-Mission Report

Crew 290 – Project MADMEN

Jan 7th – Jan 20th, 2024

Crew Members:

Commander and Health and Safety Officer: Madelyn Hoying

Executive Officer and Health and Safety Officer: Rebecca McCallin

Crew Scientist: Anja Sheppard

Green Hab Officer: Benjamin Kazimer

Crew Engineer: Anna Tretiakova

Crew Journalist: Wing Lam (Nicole) Chan

Crew Projects:

Title: Project MADMEN

Author(s): Madelyn Hoying and Rebecca McCallin, with full-crew participation

Objectives: Our mission objective is to identify and characterize microbial life via metabolic assays based on the sulfur cycle.

Current Status: Soil samples have been collected from 5 field sites with diverse geologic profiles that indicate potential for microbial activity. Measurements in the field include salinity, temperature, and ATP readings at surface level, 3-inch depth, and 6-inch depth at each site. Starting on EVA 07, pH measurements are also conducted in the field at surface and 6-inch depth. In the Science Dome, these samples are diluted and incubated in our novel microfluidic device to promote microbial growth for detection, then flushed after 24 hours to investigate through microscopy, with our first set of samples showing growth. pH measurements are taken of the soil samples in a dilution with various salts found on Mars that could also promote metabolic activity of extremophiles. The crew is targeting evaluation of at least 6 sample sites. Collected samples are compared for further analysis in the Science Dome based on depth profiles of the geology (uncovered while digging) and quality of field measurements obtained, driving collection from more than the 6 sites required for full analysis.

EVAs: 3 (EVA 04, 06, 07). One field site (EVA 04) was collected from a member of the Curtis foundation, where gypsum and sandstone were prominent under a smectite bed indicating a depositional environment with water followed by a period of dry climate. EVA 06 resulted in 2 field sites: one from a valley between two Brushy Basin members with evidence of anhydrite, to contrast with another collection site in a dried riverbed with conglomerate oyster reefs. EVA 07 saw sample collection in alternating siltstone and mudstone bands with gypsum deposits, with field pH measurements introduced into the procedure. An additional 3-4 EVAs are currently anticipated to select the best sites for further investigation.

Title: Evaluating Psychosocial Impacts of Mars Mission Architectures

Author: Madelyn Hoying

Objectives: This project seeks to compare psychosocial interactions among crew and emergency response capabilities between Mars mission architectures. Results from this single-site architecture test will be compared to previous dual-site architecture experiments developed and tested by MIT.

Current Status: The on-site investigator does not read questionnaire results while participating in the analog mission; as such, a “current status” check can only show the number of completed surveys. All participants have been submitting daily surveys, with one participant having missed one survey.

EVAs: None (although EVA inputs from other projects are valuable to the study).

Title: Ground Penetrating Radar for Martian Rovers

Author: Anja Sheppard

Objectives: This project is focused on collecting Ground Penetrating Radar (GPR) and stereoscopic camera imagery data in a Martian analog environment. Post-analog data processing will focus on using this data for machine learning applications.

Current Status: After some reconstruction and problem solving from shipping, REMI the robot is at a nominal status for data collection. REMI has collected data at 16 field sites around the MDRS Hab and at Tharsis Montes, making for over 200 GB of camera, radar, and GPS data. Anja (Crew Scientist) has been perfecting the field protocols for running REMI during EVAs, including characterizing the battery performance during cold temperatures. REMI is primarily focused on data collection, as data processing will occur post-analog.

EVAs: 4

Copyright © The Mars Society. All rights reserved. | Main Site